OpenStudy (anonymous):

there are 16 students who choose up sides to play soccer. How many different teams of 8 can be formed?

OpenStudy (anonymous):

\[16 C 8=\frac{ factoral16 }{ factorial8timesfactorrial \left( 16-8 \right) }\] =\[\frac{ factorial 16 }{ factorial 8 \times factorial 8 }\]